# MPX: Mixed Precision Training for JAX

### v2

@@ 3.2 Transformations to Cast Functions
For casting entire functions, MPX provides the transformation cast_function(func, dtype, return_dtype=None). This transformation returns a new function that first applies the specified input data type to all inputs (via cast_tree(tree, dtype)), then calls the original function, and finally casts the outputs to the specified (optional) return data type. Furthermore, MPX includes force_full_precision(func, return_dtype), which ensures computations are carried out in full precision. This feature is particularly important for operations that are prone to overflow in float16, such as sum, mean, or softmax.
@@ 3.3 Automatic Loss Scaling
(a) Full Precision Training via Equinox
(b) Half Precision Training via Equinox and MPX
Figure 1: Transformation of training pipelines via MPX. An existing training pipeline (Top figure) can be transformed, by applying @mpx.filter_grad to the forward pass. It automatically casts all inputs to half precision and calculates the gradient via dynamic loss scaling and eqx.filter_grad.
@@ 3.4 Transformations to Cast Gradient Calculation
MPX provides Equinox-compatible versions of gradient-calculation functions: filter_grad(func, scaling, has_aux=False, use_mixed_precision=True) and filter_value_and_grad(func, scaling, has_aux=False, use_mixed_precision=True). These functions compute gradients using mixed precision and loss scaling (as described in Section3.3) while automatically adjusting the loss scaling value. They act as drop-in replacements for the corresponding Equinox functions, enabling users to reuse existing training pipelines with minimal modifications (cf. Section4.2).
To use these functions, the model’s forward pass and loss computation should be combined into a single function, which is then passed to mpx.filter_value_and_grad in place of eqx.filter_value_and_grad. The transformed function then operates as follows (Figure 1):
@@ 4 Example
4.1 Model Implementation
Example 4.1 (t).
{pbox}
[] {Verbatim}[commandchars=
{},codes=] class MultiHeadAttentionBlock(eqx.Module): dense_qs: eqx.nn.Linear dense_ks: eqx.nn.Linear dense_vs: eqx.nn.Linear dense_o: eqx.nn.Linear num_heads: int layer_norm: eqx.nn.LayerNorm
def __init__(self, feature_dim, num_heads, key): self.num_heads = num_heads key, subkey = jax.random.split(key) self.dense_qs = eqx.nn.Linear( feature_dim, feature_dim, key=subkey) # same for dense_ks, dense_vs, dense_o
self.layer_norm = eqx.nn.LayerNorm(feature_dim)
def attention(q, k, v): attention_scores = q @ k.T / jnp.sqrt(q.shape[-1]) attention_scores = mpx.force_full_precision( jax.nn.softmax, attention_scores.dtype)(attention_scores, axis=-1) return attention_scores @ v
def __call__(self, inputs): inputs_after_layernorm = jax.vmap(mpx.force_full_precision( self.layer_norm, inputs.dtype))(inputs) qs = jax.vmap(self.dense_qs)(inputs_after_layernorm) qs = es.jax_einshape("n(hf)->hnf", qs, h=self.num_heads) # same for ks and vs…
outputs = jax.vmap(self.attention, in_axes=(0, 0, 0))(qs, ks, vs) outputs = es.jax_einshape("hnf->n(hf)", outputs) outputs = jax.vmap(self.dense_o)(outputs) outputs += inputs
return outputs Implementation of multi-headed self-attention via MPX. For the largest part, the model can be implemented in common toolboxes (here, Equinox (Kidger and Garcia, 2021)). Only critical function that should be implemented in full precision, like softmax or layernorm (due to mean and std calculations), have to be transformed via MPX.
For the largest part, the implementation of the model must not be changed. As MPX works with arbitrary PyTrees, every Toolbox that defines their model/parameters as PyTrees, like Flax (Heek et al., 2024) or Equinox (Kidger and Garcia, 2021) can be used in conjunction with MPX (Example 4.1). Only parts of the model that contain critical operations with a high risk of overflows, like sum or mean, must be transformed to full precision via mpx.force_full_precision.
4.2 Model Updates
@@
[] {Verbatim}[commandchars=
{},codes=] grads = eqx.filter_grad(loss)(model, batch) updates, optimizer_state = optimizer.update( grads, optimizer_state, eqx.filter(model, eqx.is_array)) model = eqx.apply_updates(model, updates)
@@
[] {Verbatim}[commandchars=
{},codes=] loss_scaling, grads_finite, grads = mpx.filter_grad(loss, loss_scaling)( model, batch) model, optimizer_state = mpx.optimizer_update( model, optimizer, optimizer_state, grads,grads_finite)
@@ 5 Evaluation
$00010001{,}$,10011001{,}1001 , 100005555101010101515151520202020Batch SizeTraining Time per Epoch (s)Full PrecisionMixed Precision
@@
002002002002004004004004006006006006008008008008001,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 8002,00020002{,}0002 , 0002,20022002{,}2002 , 20000200200200200400400400400600600600600Batch SizeTraining Time per Epoch (s)Full PrecisionMixed Precision
(b) High-Performance Cluster (4×4$\times$4 × H100)
@@
Mixed precision training via MPX reduces both RAM and processing times. For the desktop PC, the RAM is reduced by factor 1.8×1.8$\times$1.8 × and processing time by 1.7×1.7$\times$1.7 ×. The speedup is caused by reduced memory loading times as the RTX4070 streaming-multiprocessors feature no computing speedup for half precision. The cluster features a similar behavior, where processing time is reduced by up to 1.57×1.57$\times$1.57 ×.

### v3

@@ 3.2 Transformations to Cast Functions
For casting entire functions, MPX provides the transformation cast_function(func, dtype, return_dtype=None). This transformation returns a new function that first applies the specified input data type to all inputs (via cast_tree(tree, dtype)), then calls the original function, and finally casts the outputs to the specified (optional) return data type. Furthermore, MPX includes force_full_precision(func, return_dtype), which ensures computations are carried out in full precision. This feature is particularly important for operations that are prone to overflow in float16.
@@ 3.4 Transformations to Cast Gradient Calculation
MPX provides Equinox-compatible versions of gradient-calculation functions: filter_grad(func, scaling, has_aux=False, use_mixed_precision=True) and filter_value_and_grad(func, scaling, has_aux=False, use_mixed_precision=True). These functions compute gradients using mixed precision and loss scaling (as described in Section3.3) while automatically adjusting the loss scaling value. They act as drop-in replacements for the corresponding Equinox functions, enabling users to reuse existing training pipelines with minimal modifications (cf. Section4).
To use these functions, the model’s forward pass and loss computation should be combined into a single function, which is then passed to mpx.filter_value_and_grad in place of eqx.filter_value_and_grad. The transformed function then operates as follows (Figure LABEL:fig:overview):
@@ 4 Example
For the largest part, the implementation of the model must not be changed. As MPX works with arbitrary PyTrees, every Toolbox that defines their model/parameters as PyTrees, like Flax (Heek et al., 2024) or Equinox (Kidger and Garcia, 2021) can be used in conjunction with MPX. Only parts of the model that contain critical operations with a high risk of overflows, must be transformed to full precision via mpx.force_full_precision. However, many JAX and also Equinox already implement many critical functions in a numerically stable way by casting to full precision if necessary, so that only few changes are necessary.
@@
grads = eqx.filter_grad(loss)(model, batch)updates, optimizer_state = optimizer.update( grads, optimizer_state, eqx.filter(model, eqx.is_array))model = eqx.apply_updates(model, updates)
@@
loss_scaling, grads_finite, grads = mpx.filter_grad(loss, loss_scaling)( model, batch)model, optimizer_state = mpx.optimizer_update( model, optimizer, optimizer_state, grads,grads_finite)
@@ 5 Evaluation
02002004004006006008008001,0001{,}000055101015152020Batch SizeTraining Time per Epoch (s)Full PrecisionMixed Precision
@@
05005001,0001{,}0001,5001{,}5002,0002{,}0000200200400400600600Batch SizeTraining Time per Epoch (s)Full PrecisionMixed Precision
(b) High-Performance Cluster ($4\times$ H100)
@@
Mixed precision training via MPX reduces both RAM and processing times. For the desktop PC, the RAM is reduced by factor $1.8\times$ and processing time by $1.7\times$. The speedup is caused by reduced memory loading times as the RTX4070 streaming-multiprocessors feature no computing speedup for half precision. The cluster features a similar behavior, where processing time is reduced by up to $1.57\times$.